OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6y-1=6(x+5) y+5=6(x-1) y+1=6(x-5) y-5=6(x+1) Write the point-slope form of the equation passing through (5, -1) with a slope of 6y-1=6(x+5) y+5=6(x-1) y+1=6(x-5) y-5=6(x+1) @Mathematics

OpenStudy (alfie):

\[y-y_0 = m(x-x_0)\] m=6 y0 = -1 x0 = 5 Draw your conclusions.

OpenStudy (anonymous):

which one the first 1,2,3,4